[Hallc_running] [New Logentry] RC Daily Update

dhamilto at jlab.org dhamilto at jlab.org
Mon Dec 3 10:40:05 EST 2018

Logentry Text:

- Data-taking continued to go well up to 9am, at which point the beam studies started. Swing/Owl took dummy target data at 35 uA.
- The SHMS has been moved to 9.79 degrees in advance of data-taking restarting when the beam comes back around 2pm.
- The hall is at controlled access. Mark and Ethan are currently checking the HMS collimator motion controller.
- Eric traced the problem with BCM4C to the seating of cables in one of the level translators (https://logbooks.jlab.org/entry/3634399).

RC Meeting:
- Let's have a short RC meeting immediately after the Hall C meeting in A110, which is at 3pm.

Notices in effect for shift crews:
- Shift leaders, please fill out the check list once per shift.
- Shift leaders, please do a quick cross-check of the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.
- Be on the lookout for any errors when running the replay, particularly errors concerning missing reference times.
- Note the new beam positions for this energy (on the whiteboard / note on wall)
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''

Planned accelerator configuration changes:
- Beam studies / system recovery planned for Monday, Day shift -- back at 2pm.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3634411
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181203/54f5f311/attachment.html>

More information about the Hallc_running mailing list